OpenStudy (anonymous):

Solve for x. 6(2x-2)=4(x+1)

OpenStudy (anonymous):

@iGreen

OpenStudy (muzzack):

x=2

OpenStudy (anonymous):

12x-12=4x+4 12x-16=4x -16=-8x x=2